Documenta Mathematica

Band 10, 2005

Manfred Knebusch and Digen Zhang Convexity, Valuations

and Prüfer Extensions in Real Algebra  1–109

A. Vishik

On the Chow Groups of Quadratic Grassmannians  111–130

Ambrus Pál On the Torsion of the Mordell-Weil Group of the Jacobian

of Drinfeld Modular Curves  131–198

Anton Deitmar and Joachim Hilgert Cohomology of Arithmetic Groups

with Infinite Dimensional Coefficient Spaces  199–216

Peter W. Michor, David Mumford Vanishing Geodesic Distance

on Spaces of Submanifolds and Diffeomorphisms  217–245

K. Dykema, K. Jung, D. Shlyakhtenko The Microstates Free Entropy Dimension

of any DT–operator is 2  247–261

C. Cornut and V. Vatsal

CM Points and Quaternion Algebras  263–309

Robin de Jong

Arakelov Invariants of Riemann Surfaces  311–329

D. H. Jakubassa-Amundsen The Projected Single-Particle Dirac Operator

for Coulombic Potentials  331–356

Daniel Dugger, Daniel C. Isaksen

Algebraic K-Theory and Sums-of-Squares Formulas  357–366

Ekaterina Amerik Some Remarks on Morphisms Between Fano Threefolds: Erratum

cf. Documenta Math. 9, p. 471–486  367–368

Norbert Hoffmann, Ulrich Stuhler Moduli Schemes

of Generically Simple Azumaya Modules  369–389

Jean-Louis Loday and Michael R. Stein

Parametrized Braid Groups of Chevalley Groups  391–416

D. H. Jakubassa-Amundsen Localization of the Essential Spectrum

for Relativistic N-Electron Ions and Atoms  417–445

Kiran S. Kedlaya

Slope Filtrations Revisited  447–525

M. P. Dussan and M. A. Magid Time-Like Isothermic Surfaces

Associated to Grassmannian Systems  527–549

Bernhard Keller

On Triangulated Orbit Categories  551–581

Art M. Duval A Common Recursion For Laplacians of

Matroids and Shifted Simplicial Complexes  583–618

J. I. Burgos Gil, J. Kramer, U. Kühn Arithmetic Characteristic Classes

of Automorphic Vector Bundles  619–716